# Safe MPC Alignment with Human Directional Feedback

### v2

@@ Abstract
In safety-critical robot planning or control, manually specifying safety constraints or learning them from demonstrations can be challenging. In this article, we propose a certifiable alignment method for a robot to learn a safety constraint in its model predictive control (MPC) policy with human online directional feedback. To our knowledge, it is the first method to learn safety constraints from human feedback. The proposed method is based on an empirical observation: human directional feedback, when available, tends to guide the robot toward safer regions. The method only requires the direction of human feedback to update the learning hypothesis space. It is certifiable, providing an upper bound on the total number of human feedback in the case of successful learning, or declaring the hypothesis misspecification, i.e., the true implicit safety constraint cannot be found within the specified hypothesis space. We evaluated the proposed method using numerical examples and user studies in two simulation games. Additionally, we implemented and tested the proposed method on a real-world Franka robot arm performing mobile water-pouring tasks. The results demonstrate the efficacy and efficiency of our method, showing that it enables a robot to successfully learn safety constraints with a small handful (tens) of human directional corrections.
Index Terms:
Learning from Human Feedback, Constraint Inference, Model Predictive Control, Human-Robot Interaction, Cutting-Plane Method.
@@ I Introduction
In robot policy design, safety is always the top priority, especially in safety-critical applications, such as autonomous vehicles or robots operating near humans, where safety violation can lead to irreversible loss or damage. While extensive work in safe control [1] and reinforcement learning [2] focuses on designing safe policy given robot motion constraints, less attention has been paid to how to specify the safety constraints in the first place. In many applications, specifying proper safety constraints can be challenging. First, explicit safety boundary for robot behavior can be hard to model. For example, when delivering food in a crowded restaurant, a service robot must balance the tray to avoid spills; this requires constraints on robot speed, food weight distributions, and interaction with its surroundings, and all those constraint aspects can be complex to model. Second, safety perception can vary among users. In the same example of the delivery robot: some diner may feel secure with the robot staying within a foot or two when delivering items, while others may feel uncomfortable with close proximity and prefer the robot to stay at a slight distance.
@@
Figure 1: Overall illustration of the proposed safe MPC alignment. Our method online updates the safety constraint in robot safe MPC policy based on human directional feedback, such that the robot eventually aligns the safe MPC policy with human intent for accomplishing safety-critical tasks.
@@
The challenges in safety specifications lead to two questions, which this paper seeks to answer. Q1: can a robot learn and align its safety constraints with human intent in an interactive manner? Q2: can such learning process to be user-data efficient such that a robot can quickly adapt to different users? Recent work [3, 4, 5] has explored learning safety constraints from human demonstrations. However, acquiring safety-informative demonstrations could be challenging. An alternative is learning from human preference feedback [6, 7, 8, 9, 10, 11, 12], where users express preferences between trajectories to guide robot learning. Despite the success in reward alignment [6], the query trajectories can be difficult for humans to judge when they are equally suboptimal, especially in early learning stages. To address such issue, [13, 14, 15, 16, 17, 18, 19] proposed learning from human physical corrections. While promising, the recent work [20] suggests those methods are potentially sensitive to correction magnitude: improper human corrections can lead to overshooting and inefficiency of learning. Importantly, all those feedback learning methods primarily focus on learning the reward function and do not address the challenges of learning safety constraints from human feedback.
@@
This paper aims to answer the previous two questions by proposing a new paradigm for a robot to efficiently learn safety constraints in its model predictive control (MPC) policy with human feedback. We name the method safe MPC alignment. As shown in Fig. 1, a human user, observing the robot’s potentially unsafe behavior, provides feedback to the robot. Upon receiving human feedback, the robot immediately updates the safety constraints in its MPC policy and continues with the updated policy. Different from existing learning from feedback methods, our method only requires human directional corrections [20], which can be sparse in time. For example, when a human user corrects an autonomous car, he/she only needs to give the command of ‘turn left‘ or ‘turn right‘, without dictating how much the car needs to turn. The safe MPC alignment only utilizes the direction of human correction, regardless of the magnitude, for the safety constraints update. As shown in our previous work [20] or later, the type of directional feedback provides a consistently large and flexible correction space, effectively overcoming the over-shooting correction, compared to existing magnitude feedback. Importantly, the proposed alignment is able to guarantee the exponential convergence with finite rounds of human corrections, or output failure with certified hypothesis misspecification. Specifically, we highlight our main contributions below.
@@
(I) We propose the method of safe MPC alignment with human directional feedback. The method enables a robot to online learn a safe MPC policy from human directional corrections. To our knowledge, it is the first work for interactive learning of safety constraints from human feedback.
(II) We theoretically show the human-data efficiency of the safe MPC alignment: the proposed method is guaranteed to have exponential convergence with finite human corrections. Furthermore, we prove that the proposed method enables the certification of hypothesis space misspecification failure.
@@
(III) We conduct extensive user studies to validate the proposed method, including two MuJoCo-based simulation games and a real-world experiment of teaching a Franka robot arm for mobile liquid pouring. The results show the effectiveness and efficiency of our method.
@@
The remaining paper is organized as follows. Section II reviews the related work, and Section III formulates the safe MPC alignment problem. The main method is presented in Section IV. Section V gives the implementation details and presents the main results. Sections VI and VII provide the experiments and user studies. Conclusions are in Section VIII.
@@ II-A Learning from Demonstrations
Under the general theme of learning from human data, the proposed method is loosely connected to the topic of learning from demonstrations, where cost/reward functions are learned from expert (human) data using inverse reinforcement learning (IRL) and inverse optimal control (IOC) methods [21, 22]. IRL/IOC methods typically run offline: human demonstration data is first obtained, and then reward/cost learning is performed [22, 23, 24, 25, 26, 27]. As pointed in [28], acquiring optimal demonstration data can be difficult in practice. The challenge can be either attributed to the higher degrees of freedom of the robot, or the lack of prior knowledge about the desired robot behaviour to perform demonstrations with high-quality [29].
In this work, we focus on interactive learning of safety constraints. Different from learning from demonstrations that needs the exhibition of desired robot motion, the proposed method only requires a human user to incrementally improve the current robot motion by providing directional feedback, thus more accessible to human users.
@@ II-B Learning from Human Feedback
The line of this work typically focuses on learning rewards/costs from human feedback, which can be given in preference and corrections. In learning from human preferences, human preference is typically given as user’s ranking of two or multiple robot candidate trajectories [6, 7, 8, 9, 10, 11, 12], and Bradley-Terry model [30] is used to update the rewards/costs. The recent variant seeks to directly learn a policy from human preference [31].
Human feedback can also come as corrections, where human users may intervene and correct the robot at runtime. An intended robot reward/cost function can be inferred from human corrections. Specific human corrections include language [32, 33, 34, 35, 36] and physical corrections [13, 15, 18]. Since human corrections can be sparse in time, many methods [13, 14, 15, 16, 17, 18, 19] use trajectory deformation methods to interpolate sparse corrections as a continuous trajectory. The rewards/costs are updated to explain the deformed trajectories. Some strategies are further proposed to deal with the uncertainty and improve the robustness of the learning process, like updating one parameter dimension at a time [15] or formulating the reward/cost learning in the framework of Kalman Filter [14].
Our work belongs to the category of learning from human corrections. But instead of learning rewards/costs, we focus on learning safety constraints from human corrections, which has not been addressed before. Methodologically, we propose a certifiable method, which guarantees the exponential learning convergence with finite human corrections, and certifies the hypothesis misspecification if the learning hypothesis space does not contain the true human intent. Those theoretical results are missing in existing methods.
@@ II-C Learning Constraints from Demonstrations
An increasing number of work focuses on learning safety constraints from demonstrations. One vanilla idea is to formulate constraint learning as a binary classification. However, since a robot usually runs within a safe region, negative demonstration samples that violate safety are challenging to collect. To address such, [3, 37] consider the trajectories of lower cost as the unsafe samples. [38] proposes adjusting the constraints toward decreasing the safety margin of feasible policies. Safety constraints can also be learned via mixed integer programming [3, 37], pessimistic adjustment [38], and no-regret algorithm [39].
Another line of work attempts to learn safety constraints only from safe demonstration data. For example, some work exploits the optimality of demonstration with unknown constraints [40, 41, 42]. Different probabilistic formulations are also proposed to learn constraints from demonstration data, such as maximum likelihood [5, 43, 44], maximum causal entropy [45, 46], and Bayesian inference [47, 48]. In the above constraint learning, safety constraints can be represented as a set of safe states [40], the belief of potential constraints [41], Gaussian process [42], or directly as a convex hull in policy space [4].
The vast majority of the above methods formulate the constraint learning problem in the setting of reinforcement learning, and thus the learning process usually involves repetitive policy optimization. Those methods cannot be deployed in an online setting, where a robot is expected to on-the-fly update its safe policy and respond to human feedback. To fill this gap, we will develop the first method, to our best knowledge, to enable online learning of safe constraint from human feedback.
@@ III-A Safe Model Predictive Control
We consider the following safe model predictive control (Safe MPC) for a robotic system, where the safety constraint is learnable and parameterized by $\boldsymbol{\theta}\in\mathbb{R}^{r}$.
$\min_{\boldsymbol{u}_{0:T-1}}&\quad J(\boldsymbol{u}_{0:T-1};\boldsymbol{x}_{0})=\sum_{t=0}^{T-1}c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})+h(\boldsymbol{x}_{T})\\ \text{subject to}&\quad\boldsymbol{x}_{t+1}=\boldsymbol{f}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})\\ &\quad{g}_{\boldsymbol{\theta}}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\leq 0\end{aligned}.$ (1)
Here, $\boldsymbol{x}_{t}\in\mathbb{R}^{n}$ and $\boldsymbol{u}_{t}\in\mathbb{R}^{m}$ are robot state and action, respectively; $t=0,1,2,...,T$ is the time step with prediction horizon $T$; $\boldsymbol{x}_{0:T}=(\boldsymbol{x}_{0},\boldsymbol{x}_{1},...,\boldsymbol{x}_{T})$ and $\boldsymbol{u}_{0:T-1}=(\boldsymbol{u}_{0},\boldsymbol{u}_{1},...,\boldsymbol{u}_{T-1})$; $J(\cdot)$ is the given cost function including the running cost ${c}(\cdot)$ and final cost $h(\cdot)$, representing the given task; $\boldsymbol{f}(\cdot,\cdot)$ is the system dynamics; ${g}_{\boldsymbol{\theta}}(\cdot)$ is the safety constraint, which is parameterized by $\boldsymbol{\theta}\in\mathbb{R}^{r}$. Given starting state $\boldsymbol{x}_{0}$ and constraint parameter $\boldsymbol{\theta}$, the trajectory solution to (1), denoted as $\boldsymbol{\xi}_{\boldsymbol{\theta}}=\{\boldsymbol{u}^{\boldsymbol{\theta}}_{0:T-1},\boldsymbol{x}^{\boldsymbol{\theta}}_{0:T}\}$, is an implicit function of $\boldsymbol{\theta}$.
In a receding horizon implementation, $\boldsymbol{x}_{0}$ is set as the robot’s current state (estimation) $\boldsymbol{{x}}_{k}$, where $k=0,1,2,...,$ is the policy rollout time step; and only the first action $\boldsymbol{u}^{\boldsymbol{\theta}}_{0}$ in $\boldsymbol{{u}}_{0:T}^{\boldsymbol{\theta}}$ is applied to the robot to move the robot to the next state $\boldsymbol{x}_{k+1}$. This leads to a closed-loop Safe MPC policy for the robot, denoted as $\boldsymbol{\pi}_{\boldsymbol{\theta}}$, mapping from robot current state $\boldsymbol{x}_{k}$ to its action $\boldsymbol{u}^{\boldsymbol{\theta}}_{0}$.
@@ III-B Safe Asymptotic Approximation of Safe MPC
Here, the above defined penalty $B(\boldsymbol{\xi},\boldsymbol{\theta})$ includes the original control cost in (1) and safety constraint $g_{\boldsymbol{\theta}}(\boldsymbol{\xi})$ via logarithmic barrier transformation, $\gamma>0$ is named the barrier parameter, controlling the weight of safety penalty $-\ln(-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}))$ relative the control cost. With $B(\boldsymbol{\xi},\boldsymbol{\theta})$, one can establish the following penalty-based MPC formulation
@@
We denote the solution to the above penalty-based MPC (3) as $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}=\{\boldsymbol{u}_{0:T-1}^{\boldsymbol{\theta},\gamma},\boldsymbol{x}_{0:T-1}^{\boldsymbol{\theta},\gamma}\}$. The following result is from our prior work [49], stating that the penalty-based MPC (3) is an asymptotic and safe approximation to the Safe MPC (1).
@@
(Theorem 2 in [49]) If the original Safe MPC (1) satisfies certain mild conditions (i.e., local second-order condition and strict complementarity), then for any small $\gamma>0$, the penalty-based MPC (3) has a (local) solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, and $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is safe with respect to the original constraint, i.e., $g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})<0$, and $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ satisifies the first-order condition of (3). Furthermore, $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is asymptotically approximate $\xi_{\boldsymbol{\theta}}$: $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\rightarrow\boldsymbol{\xi}_{\boldsymbol{\theta}}$ as $\gamma\rightarrow 0$.
The above lemma asserts that solving the Safe MPC (1) can be replaced by solving penalty-based MPC (3). The solution approximation is controlled by $\gamma$ for arbitrary accuracy, as $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\rightarrow\boldsymbol{\xi}_{\boldsymbol{\theta}}$ as $\gamma\rightarrow 0$. A good approximation is achieved by smaller $\gamma$. on the contrary, if $\gamma$ is larger, penalty-based MPC (3) will lead to more safety-conservative solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$. In the development below, we fix the parameter $\gamma$ in (3). We denote the penalty-based MPC policy in (3) as $\boldsymbol{\pi}^{\gamma}_{\theta}$.

### v3

@@ Abstract
In safety-critical robot planning or control, manually specifying safety constraints or learning them from demonstrations can be challenging. In this article, we propose a certifiable alignment method for a robot to learn a safety constraint in its model predictive control (MPC) policy from human online directional feedback. To our knowledge, it is the first method to learn safety constraints from human feedback. The proposed method is based on an empirical observation: human directional feedback, when available, tends to guide the robot toward safer regions. The method only requires the direction of human feedback to update the learning hypothesis space. It is certifiable, providing an upper bound on the total number of human feedback in the case of successful learning, or declaring the hypothesis misspecification, i.e., the true safety constraint cannot be found within the specified hypothesis space. We evaluated the proposed method in numerical examples and user studies with two simulation games. Additionally, we tested the proposed method on a real-world Franka robot arm performing mobile water-pouring tasks. The results demonstrate the efficacy and efficiency of our method, showing that it enables a robot to successfully learn safety constraints with a small handful (tens) of human directional corrections.
@@ I Introduction
In robot policy design, safety is always the top priority, especially in safety-critical applications, such as robots operating near humans, where safety violations can lead to irreversible losses. While extensive work in safe control [wabersich2023data] and reinforcement learning [gu2022review] focuses on deducing robot policy given safety constraints, less attention has been paid to how to specify the safety constraints in the first place. In many applications, specifying proper safety constraints can be challenging. First, explicit boundary for safe robot behavior can be hard to model. For example, when delivering food in a crowded restaurant, a service robot must balance a tray to avoid spills; this requires constraints on robot speed, food weight distributions, and interaction with its surroundings, all of which can be complex to model. Second, user’s perception of safety may vary. In the same example, some diner may feel secure with the robot staying within a foot when delivering, while others may feel uncomfortable with such proximity.
@@
Figure 1: Illustration of the proposed safe MPC alignment. Our method online updates the safety constraint using human directional feedback, such that the robot eventually aligns the safe MPC policy with human intent for the safety-critical task.
@@
The challenges in safety specification lead to two questions. Q1: can a robot learn safety constraints to align with human intent in an interactive manner? Q2: can such a learning process be user-data efficient? Recent work [chou2020learning_1, lindner2023learning, scobee2019maximum] explored learning safety constraints from human demonstrations. However, acquiring safety-informative demonstrations could be difficult. Learning from human preference [christiano2017deep, Sadigh2017ActivePL, ibarz2018reward, pmlr-v139-lee21i, hejna2023few, 10160439, pmlr-v164-myers22a] offers a new paradigm, where a user’s preference of one trajectory over the other is used for robot learning. The preference, however, can be difficult for a human to make, especially in early learning stages. [bajcsy2017learning, losey2018including, bajcsy2018learning, zhang2019learning, li2021learning, losey2022physical, mehta2023unified] proposed learning from human physical corrections. While promising, [jin2022learning] suggests those methods are potentially sensitive to correction magnitude: corrections can lead to overshooting and inefficiency of learning. All those feedback-driven learning methods primarily focus on learning rewards instead of safety constraints.
@@
This paper aims to answer the previous two questions by proposing a new paradigm for a robot to efficiently learn safety constraints in its model predictive control (MPC) policy from human feedback. We name the method safe MPC alignment. As in Fig. 1, a user, observing the robot’s potentially unsafe behavior, provides online corrections to the robot. Upon receiving human corrections, the robot immediately updates the safety constraints in its MPC policy and continues with the updated policy. Our method only requires/utilizes direction of a human correction [jin2022learning]. For example, when a human user corrects an autonomous car, he/she only needs to give the command of ‘turn left’ or ‘turn right’, without dictating how much the car needs to turn. Directional corrections permit a consistently large and flexible correction space. Importantly, the proposed method guarantees the exponential convergence with finite rounds of human corrections, or outputs failure with certified hypothesis misspecification. Specifically, we highlight our main contributions below.
@@
(I) We propose the method of safe MPC alignment from human directional correction. It enables a robot to learn a safe MPC policy from online human directional corrections. To our knowledge, it is the first work for interactive learning of safety constraints from human feedback.
(II) We show that safe MPC alignment is certifiable: it guarantees exponential convergence with a finite number of human corrections and enables verification when the true safety constraint falls outside the specified hypothesis space.
@@
(III) Extensive user studies have been conducted to validate the proposed method, including a real-world experiment on the Franka robot arm. The results show the effectiveness and efficiency of our method.
@@
The remaining paper is organized as follows. Section II reviews the related work, and Section III formulates the safe MPC alignment problem. The main method is presented in Section IV. Section V gives the implementation details and the main results. Sections VI-VIII provide the experiments and user studies. Discussions and Conclusions are in Section IX.
@@ II-A Learning from Demonstrations
Under the general theme of learning from human inputs, the proposed method falls into the category of learning from demonstrations, where costs/rewards are learned from expert data using inverse reinforcement learning (IRL) [ng2000algorithms]. IRL methods typically run offline: human demonstrations are first collected, and then reward/cost is learned [ng2000algorithms, abbeel2004apprenticeship, ratliff2006maximum, ziebart2008maximum, mombaur2010human, englert2017inverse]. As pointed by [ravichandar2020recent], acquiring demonstration data can be difficult in practice, due to either the high degrees of freedom of a robot, or the lack of knowledge about the desired robot behavior [jain2015learning].
In this work, we focus on interactive learning of safety constraints. Different from learning from demonstrations, the proposed method only requires a user to incrementally improve the current robot motion by providing directional feedback.
@@ II-B Learning from Human Feedback
This line typically focuses on learning rewards/costs from human feedback in forms of preferences and corrections. Human preference is typically given as user’s ranking of two or multiple robot trajectories [christiano2017deep, Sadigh2017ActivePL, ibarz2018reward, pmlr-v139-lee21i, hejna2023few, 10160439, pmlr-v164-myers22a], which is modeled using the Bradley-Terry formulation [bradley1952rank]. Human corrections require a user to intervene and correct the robot at runtime, from which a intended reward/cost can be inferred. Human corrections can include language command [karamcheti2022shared, cui2023no, srivastava2023generating, zha23droc, shi2024yell] and physical corrections [bajcsy2017learning, bajcsy2018learning, losey2022physical]. Since human corrections can be sparse in time, many methods [bajcsy2017learning, losey2018including, bajcsy2018learning, zhang2019learning, li2021learning, losey2022physical, mehta2023unified] use trajectory deformation methods to interpolate sparse corrections as a continuous trajectory. Some strategies are further proposed to deal with learning robustness and uncertainty, such as updating one feature at a time [bajcsy2018learning] or formulating the problem in the Kalman Filter framework [losey2018including].
Our work belongs to the category of learning from human corrections. But instead of learning rewards/costs, we focus on learning safety constraints from human corrections, which has not been addressed before. Methodologically, we propose a certifiable method, which guarantees the learning convergence with finite human corrections, and certifies the hypothesis misspecification. Those results are missing in existing methods.
@@ II-C Learning Constraints from Demonstrations
An increasing number of work focuses on learning safety constraints from demonstrations. One straightforward idea is to cast the problem as a binary classification. [chou2020learning_1, chou2020learning_2] consider the trajectories of lower cost as the unsafe samples. [gaurav2023learning] proposes adjusting the constraints toward decreasing the safety margin of feasible policies. Safety constraints can also be learned via mixed integer programming [chou2020learning_1, chou2020learning_2], pessimistic adjustment [gaurav2023learning], and no-regret algorithm [kim2024learning]. Those methods typically require collecting negative (unsafe) samples, which can be challenging in real robot systems. Another line of work attempts to learn safety constraints only from safe demonstration. For example, [chou2020learning_3, chou2021uncertainty, chou2022gaussian] infer the safety constraints by exploiting the optimality conditions of demonstrations. Different probabilistic formulations are also used, such as maximum likelihood [scobee2019maximum, stocking2022maximum, malik2021inverse], causal entropy [mcpherson2021maximum, baert2023maximum], and Bayesian inference [park2020inferring, papadimitriou2022bayesian]. Among them, safety constraints can be represented as a set of safe states [chou2020learning_3], the belief of constraints [chou2021uncertainty], Gaussian process [chou2022gaussian], or a convex hull in policy space [lindner2023learning].
Using demonstrations, the vast majority of the above methods formulate the constraint learning in offline settings, and the learning process usually involves repetitive policy optimization. Those methods can be difficult to be deployed online, where a robot is expected to quickly update its safety estimate and respond to human feedback. To fill this gap, we will develop the first method, to our best knowledge, to enable online learning of safe constraints from human feedback.
@@ III-A Safe Model Predictive Control
Consider a robot equipped with a safe model predictive control (Safe MPC) policy defined below, where the safety constraint is learnable and parameterized by $\boldsymbol{\theta}\in\mathbb{R}^{r}$.
$\begin{aligned} \min_{\boldsymbol{u}_{0:T-1}}&\quad J(\boldsymbol{u}_{0:T-1};\boldsymbol{x}_{0})=\sum\nolimits_{t=0}^{T-1}c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})+h(\boldsymbol{x}_{T})\\ \text{subject to}&\quad\boldsymbol{x}_{t+1}=\boldsymbol{f}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})\quad\text{given}\quad\boldsymbol{x}_{0}\\ &\quad{g}_{\boldsymbol{\theta}}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\leq 0\end{aligned}.$ (1)
Here, $\boldsymbol{x}_{t}\in\mathbb{R}^{n}$ and $\boldsymbol{u}_{t}\in\mathbb{R}^{m}$ are robot state and action, respectively; t=0,1,2,,Tt=0,1,2,...,T is the time step over a short prediction horizon $T$, starting from a given $\boldsymbol{x}_{0}$. Define $\boldsymbol{x}_{0:T}=(\boldsymbol{x}_{0},\boldsymbol{x}_{1},...,\boldsymbol{x}_{T})$ and $\boldsymbol{u}_{0:T-1}=(\boldsymbol{u}_{0},\boldsymbol{u}_{1},...,\boldsymbol{u}_{T-1})$; $J(\cdot)$ is the given cost function including the running cost ${c}(\cdot)$ and final cost $h(\cdot)$; $\boldsymbol{f}(\cdot,\cdot)$ is the system dynamics; ${g}_{\boldsymbol{\theta}}(\cdot)$ is the safety constraint parameterized by $\boldsymbol{\theta}\in\mathbb{R}^{r}$. Given starting state $\boldsymbol{x}_{0}$ and parameter $\boldsymbol{\theta}$, the optimal plan (solution) to (1), denoted as $\boldsymbol{\xi}_{\boldsymbol{\theta}}=\{\boldsymbol{u}^{\boldsymbol{\theta}}_{0:T-1},\boldsymbol{x}^{\boldsymbol{\theta}}_{0:T}\}$, is an implicit function of $\boldsymbol{\theta}$.
In a receding horizon implementation, $\boldsymbol{x}_{0}$ is set as the real robot’s current state (estimation) $\boldsymbol{{x}}_{k}^{\text{real}}$, where k=0,1,2,,k=0,1,2,..., is the global rollout time step of the Safe MPC policy. Only the first action $\boldsymbol{u}^{\boldsymbol{\theta}}_{0}$ in each $\boldsymbol{\xi}_{\boldsymbol{\theta}}$ is applied to the robot to move the robot to the next state $\boldsymbol{x}^{\text{real}}_{k+1}$. This process repeats at each rollout time step $k=0,1,$.... Such a receding horizon implementation leads to a closed-loop policy, denoted as $\boldsymbol{\pi}_{\boldsymbol{\theta}}$, mapping from robot current state $\boldsymbol{x}_{k}^{\text{real}}$ to its action $\boldsymbol{u}^{\boldsymbol{\theta}}_{0}$.
@@ III-B Safe Asymptotic Approximation of Safe MPC
Here, $B(\boldsymbol{\xi},\boldsymbol{\theta})$ includes the cost $J(\boldsymbol{\xi})$ and safety constraint $g_{\boldsymbol{\theta}}(\boldsymbol{\xi})$ in a logarithmic barrier transformation. $\gamma>0$ is the barrier parameter, controlling the weight of safety penalty $\ln(-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}))$ relative the cost. With $B(\boldsymbol{\xi},\boldsymbol{\theta})$, one can establish the following penalty-based MPC
@@
We denote the optimal plan to (3) as $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}=\{\boldsymbol{u}_{0:T-1}^{\boldsymbol{\theta},\gamma},\boldsymbol{x}_{0:T-1}^{\boldsymbol{\theta},\gamma}\}$. The following result from [jin2021safe] states the penalty-based MPC (3) is an asymptotic, safe approximation to Safe MPC (1).
@@
(Theorem 2 in [jin2021safe]) If the Safe MPC (1) satisfies certain mild conditions (i.e., local second-order condition and strict complementarity), then for any small $\gamma>0$, the penalty-based MPC (3) has a (local) solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, and $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is safe with respect to the original constraint, i.e., $g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})<0$, and $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ satisifies the first-order condition of (3). Furthermore, $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is asymptotically approximate $\xi_{\boldsymbol{\theta}}$: $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\rightarrow\boldsymbol{\xi}_{\boldsymbol{\theta}}$ as $\gamma\rightarrow 0$.
The above lemma asserts that solving the Safe MPC (1) can be replaced by solving penalty-based MPC (3). The solution approximation is controlled by $\gamma$ for arbitrary accuracy: $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\rightarrow\boldsymbol{\xi}_{\boldsymbol{\theta}}$ as $\gamma\rightarrow 0$. A good approximation is achieved by a smaller $\gamma$. On the contrary, if $\gamma$ is larger, penalty-based MPC will lead to more safety-conservative solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$. In the development below, we fix the parameter $\gamma$ in (3). We denote the penalty-based MPC policy in (3) as $\boldsymbol{\pi}^{\gamma}_{\theta}$.